OpenStudy (anonymous):

Find the equation of the parabola with focus (-3,1), and vertex (-4,1)

OpenStudy (anonymous):

answer or method?

OpenStudy (anonymous):

How to do it o;

OpenStudy (anonymous):

Tysm C:

OpenStudy (anonymous):

actually this form is a little misleading so lets do it the right way vertex is \((-4,1)\) and we know it opens to the right (because the focus is one unit right, namely \((-3,1)\) so we start with \[(y-1)^2=a(x+4)\] and then we need to find \(a\)

OpenStudy (anonymous):

in this case you have gone one unit to the right from the vertex to the focus, so \(a=4\times 1=4\) and so you should really write \[(y-1)^2=4(x+4)\]

OpenStudy (anonymous):

so if your parabola opens right with vertex at \((h,k)\) you know it looks like \[(y-k)^2=a(x-h)\] and find \(a\) by seeing how many units you have gone to the right to get the focus, and multiply by 4

OpenStudy (anonymous):

Much better xD Thanks again